I am lost on this equation solution. I got the answer from the answer sheet but still can't figure out how to get there.
x/3 + y/6 = 1
x/2 - y/4 = 0
answer is 3/2, 3
How?

Answer by richard1234(7193)

Multiply first equation by 6 and second equation by 4 (this clears the fractions)

2x + y = 6
2x - y = 0

Add both equations together to clear the y terms

4x = 6 --> x = 3/2

Replace x into either equation to obtain y = 3.